OpenStudy (anonymous):

I will give you a medal and be your fan What is 7*2 3/7 ? A. 7 6/7 B. 14 3/7 C. 14 6/7 D. 17 I think it's C

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Its 17.

OpenStudy (anonymous):

What math is this?

OpenStudy (anonymous):

It's pre-algebra and @xxccjxx can you explain the answer?

OpenStudy (anonymous):

ok i get it now, you forgot to properly write the equation 7*(2 3/7). Just multiply it in a calculator and you get 17

OpenStudy (anonymous):

You multiply 7 by 2 3/7. So you if you multiply 7 by 2, yout get 14. and if you multiply 7 by 3/7 you get 3 so add 14 and 3 together and you get 17.

OpenStudy (anonymous):

xxcc you lost me after you said multiply by 2 and get 14..Everything after I don't understand and @Fransisco_w no it's not suppose tp be in parentheses

OpenStudy (anonymous):

Multipy 7 x 2

OpenStudy (anonymous):

Yout get 14

OpenStudy (anonymous):

multiply them separately and add them at the end

OpenStudy (anonymous):

multipy 7 by 3/7 yout get 3

OpenStudy (whpalmer4):

Yes, what @Fransisco_w wrote is correct. You can do this at least two ways: 1) convert to an improper fraction \[7*2\frac{3}{7} = 7*\frac{2*7+3}{7}= 7*\frac{17}{7} = \cancel{7}\frac{17}{\cancel{7}} = 17\] 2) use distributive property \[7*2\frac{3}{7} = 7*(2+\frac{3}7) = 7*2 + 7*\frac{3}{7} = 14 + 3 = 17\]

OpenStudy (anonymous):

Add them together and you 17

OpenStudy (anonymous):

Exactly what i said @whpalmer4

OpenStudy (anonymous):

Ohhh ok I understand it a little more now

OpenStudy (anonymous):

Thanks everyone